Mona and her grandfathers age add up to 82. Five years ago, her grandfather was 7 times as old as Mona was then. Find their ages
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

Mona and her grandfathers age add up to 82.
M + G = 82
-------
Five years ago, her grandfather was 7 times as old as Mona was then. Find their ages
----
G-5 = 7*(M-5)
etc
====================
I'm amazed at how many people don't know how old they are.

Answer by greenestamps(13200)



5 years ago their ages added to 72.

Divide 72 into two parts in the ratio 7:1 to find their ages 5 years ago.

Add 5 to each of those ages to find their current ages.